Problem 3.108 (Chapter 3)

1) If an automobile travels 245 mi with a gas mileage of 20.5 mi/gal , how many kilograms of CO2 are produced? Assume that the gasoline is composed of octane, C8H18(l), whose density is 0.69 g/mL.

2) Repeat the calculation for a truck that has a gas mileage of 7 mi/gal mi/gal.
1
Expert's answer
2016-10-09T07:49:03-0400
1) For 245 mi you need 245 mi/ 20.5 mi/gal = 12.0 gal of gasoline. 1 gal is some 4.546 lit. Thus, you need 54.6 liters of gasoline. It is some 54.6 l * 690 g/l = 37.7 kg. Octane burns according to the following equation 2C8H18 + 25O2 = 16CO2 + 18H2O. 37.7 kg of octane-gasoline is 331 mol of octane or 2650 mol of CO2 which mass is 11.7 kg of pure CO2.

2) For truck with 7 mi/gal we need 35 gal of gasoline. It is 159 liters or 159 l * 690 g/l = 110 kg. Thus, 110 kg of octane-gasoline is 965 mol of gasoline or 7720 mol of CO2 which mass is 340 kg.